# MindVote: When AI Meets the Wild West of Social Media Opinion

### v3

@@ Abstract
Large Language Models (LLMs) are increasingly used as scalable tools for pilot testing, predicting public opinion distributions before deploying costly surveys. To serve as effective pilot testing tools, the performance of these LLMs is typically benchmarked against their ability to reproduce the outcomes of past structured surveys. This evaluation paradigm, however, is misaligned with the dynamic, context-rich social media environments where public opinion is increasingly formed and expressed. By design, surveys strip away the social, cultural, and temporal context that shapes public opinion, and LLM benchmarks built on this paradigm inherit these critical limitations. To bridge this gap, we introduce MindVote, the first benchmark for public opinion distribution prediction grounded in authentic social media discourse. MindVote is constructed from 3,918 naturalistic polls sourced from Reddit and Weibo, spanning 23 topics and enriched with detailed annotations for platform, topical, and temporal context. Using this benchmark, we conduct a comprehensive evaluation of 15 LLMs. MindVote provides a robust, ecologically valid framework to move beyond survey-based evaluations and advance the development of more socially intelligent AI systems.
Keywords Opinion Prediction  $\cdot$ Large Language Models  $\cdot$ Social Media  $\cdot$ Cross-Cultural Evaluation  $\cdot$ LLM Benchmark
@@ 1 Introduction
A core application for Large Language Models (LLMs) in understanding public opinion is serving as a rapid, scalable tool for pilot testing—predicting how a population will respond to a question before deploying costly, large-scale surveys rothschild2024opportunities ; bisbee2024synthetic ; suh2025language ; sinacola2025llms ; cao2025specializing ; anthis2025llm ; binz2025should . This capability is crucial across a wide spectrum of domains. For example, LLMs can be used to gauge reactions to new entertainment content and technological innovations, understand career aspirations, navigate complex social issues, and track evolving lifestyle trends argyle2023out ; qian2025large ; chen2025large . However, the central goal is not only forecasting the majority view, but also accurately predicting the entire public opinion distribution across a range of choices because the shape of the distribution reveals crucial information about societal consensus, polarization, and the intensity of dissent, which is lost in a single majority statistic zhou-etal-2022-distributed ; moon-etal-2024-virtual ; meister-etal-2025-benchmarking ; cao-etal-2025-specializing .
@@
The prevailing paradigm for public opinion distribution prediction relies on traditional structured surveys krogstad2016pew ; santurkar2023whose ; elkjaer2024estimating ; suh2025language . This methodology is misaligned with the fast-paced, context-rich digital environments where opinions are now actively formed and expressed, such as social media platforms (boutyline2021cultural, ; riva2022social, ). While surveys include basic demographic questions—such as age, gender, or education level—these static variables are poor proxies for the dynamic, situated nature of social context zaller1992simple . Demographics may indicate who a person is in categorical terms, but they fail to reflect how opinions are shaped by lived experiences, group affiliations, ongoing conversations, or culturally specific reasoning patterns turner1979social ; kahan2011cultural ; markus2014culture ; mcintosh2019exploring .
@@
Consequently, LLM benchmarks built on this flawed paradigm, such as OpinionQA and SubPop (santurkar2023whose, ; suh2025language, ), inherit these critical limitations. This reveals a dangerous gap in the responsible deployment of LLMs. As these models are increasingly used to act as a proxy for human populations, their accuracy becomes paramount. A proper evaluation framework must identify when models fail at context-rich, real-world tasks. Without this, we risk promoting systems that misrepresent the groups they are meant to simulate. This misrepresentation has severe downstream consequences, potentially compromising strategic decisions across critical domains like public policy, marketing, technology development, and community management (qi2024impact, ; radivojevic2024llms, ; qu2024performance, ; panickssery2024llm, ; repec:arx:papers:2502.03158, ; wang2025large, ). This reliance on traditional survey-based evaluation creates three critical gaps in our ability to assess an LLM’s true social understanding:
@@
The Gap about Topic Imbalance. Real-world online discourse is thematically diverse tran-ostendorf-2016-characterizing ; song2023understanding ; entertainment, for instance, drives 70% of traffic on Weibo (chinamarketingcorp2024weibo, ), and Reddit’s largest communities focus on gaming and technology (proferes2021studying, ). Yet, current survey-based benchmarks are heavily skewed towards formal, institutional topics like politics, which represent only a fraction of naturalistic opinion expression (santurkar2023whose, ; zhao-etal-2024-worldvaluesbench, ). This mismatch means we cannot evaluate a model’s ability to adapt its reasoning from the formal discourse of a political survey to the vernacular social media discourse karjus2024evolving ; reveilhac2024augmenting .
@@
The Gap about Cultural Homogeneity. Existing cross-cultural survey benchmarks often exhibit cultural homogeneity, creating test conditions on Western-centric questions even with other languages translation (haerpfer2022world, ; durmus2024towards, ; zhao-etal-2024-worldvaluesbench, ). A model might perform well on a translated question, but we are left unable to determine if this success stems from genuine cultural understanding or a superficial linguistic competence singh-etal-2025-global .
@@
The Gap about Social Context Poverty. Opinions on social media are not noise but actual signals for prediction lopez2021framing ; they are primed by platform norms (e.g., Reddit’s anonymity vs. Weibo’s public-facing profiles (brown2018reddit, ; zhu2024privacy, )), temporal events banducci2015surveys , and community-specific discourse (tan2016winning, ; qiao2023does, ). Survey-based benchmarks systematically remove this contextual priming to achieve standardization (tourangeau2003context, ), preventing a true evaluation of a model’s ability to leverage these critical predictive cues 10.24963/ijcai.2023/749 .
@@
To overcome these limitations, we extend the opinion distribution prediction scope of survey to a source of direct public opinion formation scarano2024analyzing : naturalistic social media polls. These polls directly address the three critical gaps, providing a robust foundation for testing the very abilities that current benchmarks obscure. The comparison between existing survey-based benchmarks and our benchmark is in Figure 1.
@@
Our Contributions. We introduce MindVote, the first benchmark for public opinion distribution prediction grounded in authentic social media discourse. We advance the field with the following innovations:
  • We construct and release MindVote, a benchmark of 3,918 authentic polls from two platforms (Reddit and Weibo) in their native English and Chinese. It spans 5 major topics and 23 sub-topics (shown in Figure 2 and is enriched with detailed social context annotations served as context priming.
  • We establish performance baselines for leading LLMs, identifying the best-performing models while also uncovering that models specifically fine-tuned on traditional surveys reveal a critical “survey-based specialization pitfall."
  • We provide strong evidence that enhancing a model’s capacity for in-context social reasoning is a more robust and effective path forward than fine-tuning on context-stripped data, offering a new direction for developing more socially intelligent systems.
@@ 2 Related Work
Opinion Distribution Prediction Benchmarks. U.S.-focused benchmarks like OpinionQA (santurkar2023whose, ) and SubPop (suh2025language, ) concentrate on political topics, while international efforts like GlobalOpinionQA (durmus2024towards, ) is Western-centric frameworks and WorldValuesBench (zhao-etal-2024-worldvaluesbench, ) is framed as multi-cultural value prediction. cao2025specializing introduces three datasets (two English, one Chinese) specifically for group-level distribution prediction. These approaches contain only demographic-value pairings, structured survey format, or Western-centric lens that abstract away the naturalistic cultural contexts and context priming that are essential for authentic opinion distribution prediction.
@@ 3.1 Dataset Construction
Multi-Platform Data Collection. We collected 7,648 polls across three platforms spanning 2019-2025. Weibo dataset contributed 3,757 polls (2,026 from existing datasets lu-etal-2021-engage from 2019 to 2021, 1,734 newly crawled) and are anonymized for users’ personal identifiable information, capturing Chinese social media dynamics across pandemic and post-pandemic periods. Reddit provided 3,891 polls from diverse subreddits including r/poll during 2021-2025.
@@
    Content Quality Filtering: This filtering begins with duplicate content removal using a MinHash algorithm to eliminate items with greater than 95% overlap (shrivastava2014defense, ), effectively targeting redundant trending topics, reposted polls, and cross-platform duplicates. Subsequently, automated toxicity screening was conducted using the Google Perspective API, with polls retaining a toxicity score below 0.4 being retained to filter out hate speech and overly controversial subjects (lees2022new, ).
@@
    Social Relevance Validation: Human verification of voting patterns (removing polls $\leq$ 100 votes for social relevance) and verification of poll and vote existence in all three platforms to ensure authentic social engagement and meaningful community participation.
@@
We manually annotate each poll includes rich metadata for social context enrichment: general platform context (e.g., user statistics and user behaviors), topical context within that platform (topic-specific discourse patterns), and temporal context for each poll (social events and backgrounds related to the poll when created).
Cross-Cultural Augmentation. We created a parallel bilingual corpus for all three platforms through machine translation with rigorous quality control to demonstrate both linguistic and cultural effects: 10% back-translation validation (BLEU $>$ 35 (papineni2002bleu, )), 5% native speaker rewriting, and expert review (Fleiss’ $\kappa$ = 0.51). The total number of polls becomes 3,918 after augmentation.
Final Dataset Composition. The final MindVote dataset is composed of 3,918 polls, with 2,158 sourced from Reddit and 1,760 from Weibo. Each poll is enriched with a comprehensive set of metadata, including its creation time, total vote count, and three layers of social context: general platform context, topical context within that platform, and the specific temporal context at the time the poll was created. To ensure broad accessibility and ease of use, the entire dataset is provided in both CSV and JSON formats. Table 1 shows one example with its simplified metadata keywords for illustration.
@@
Temporal Context AI-driven layoff spike fuels job insecurity.
@@ 3.2 Experiment Design
Models. We evaluate 15 leading LLMs across closed-source, open-source and specalization categories: Claude-3.7-Sonnet-02-24 (claude-3.7-thinking, ), GPT-4o-2024-11-20 (openai-4o, ), GPT-4.1-04-14 (openai-4.1, ), Gemini-2.5-Pro-05-06 (google-gemini-2-5-pro, ), o3-medium-04-16 (openai-o3-o4mini-system-card, ), Deepseek-R1-05-28 (2501.12948, ), Qwen2.5-32B-Instruct (qwen2.5, ), Gemma-2-9B-it (team2024gemma, ), Mistral-7B-Instruct-v0.1 chaplot2023albert , Llama-2-13B-Chat touvron2023llama , Llama-3-70B-Instruct (grattafiori2024llama, ), and Llama-4-Maverick-17B-128E (llama4, ). We also evaluate three opinion distribution prediction specialization models from suh2025language . These models are fine-tuned using LoRA hu2022lora on base models: Llama-3-70B-Base grattafiori2024llama , Llama-2-13B-Base touvron2023llama , and Mistral-7B-v0.1 chaplot2023albert which we named them respectively as: SubPop-Llama-3-70B, SubPop-Llama-2-13B, SubPop-Mistral-7B.
@@
Pipeline. All primary evaluations use a greedy decoding strategy (temperature=0) with default hyperparameter settings under zero-shot with context priming. For the specialization fine-tuned models, we adapt those models into our pipeline by first loading the respective pretrained base model and then applying the publicly available LoRA weight checkpoints provided by suh2025language .
@@
Prompt. To ensure consistent and machine-readable outputs, we employ a structured prompting strategy where the model is given a JSON object containing the poll and its context. The template instructs the model to assume the role of a “opinion distribution prediction expert analyzing voting patterns and social dynamics." The prompt includes the poll question and is enriched with social context metadata served as context priming, with instruction for step-by-step reasoning. The model’s task is to return a JSON object with a schema identical to the input, replacing placeholder fields with its numeric predictions for the voting distribution. The complete and structured template is detailed in Appendix LABEL:appendix:prompt.
@@
Evaluation Metrics. We adopt four distinct metrics to provide a comprehensive evaluation. Our primary metric is 1 - Wasserstein Distance (1-Wass.) (santurkar2023whose, ; meister-etal-2025-benchmarking, ; suh2025language, ). The Wasserstein Distance measures the minimum cost for transforming one distribution into another, crucially accounting for semantic similarity between answer choices by treating them as points in a metric space. To complement this, we also report Spearman’s Rank Correlation Coefficient ($\rho$) (zhou-etal-2022-distributed, ; pavlopoulos2023distance, ), a non-parametric measure of how well the predicted ranking of options matches the true ranking of vote shares; 1 - KL Divergence (meister-etal-2025-benchmarking, ; nguyen2025task, ), which quantifies the information loss when using the model’s predicted distribution to approximate the ground truth; and One-hot Accuracy (zhou-etal-2022-distributed, ; santurkar2023whose, ; suh2025language, ), which provides a strict measure of whether the single most likely predicted answer is correct.
@@
Evaluation Boundary. We include upper bounds and lower bounds for comparisons following suh2025language . The upper bound is established by sampling subsets of the original results, calculating the four metrics between subsampled and original distributions, and performing bootstrapping to obtain a robust estimate that captures the intrinsic variance arising from the respondent sampling process in opinion. The uniform distribution lower bound establishes a performance floor equivalent to random chance.

### v5

@@ Abstract
Large Language Models (LLMs) are increasingly used as scalable tools for pilot testing, predicting public opinion distributions before deploying costly surveys. However, the prevailing paradigm for evaluating these models relies on traditional structured surveys—a methodology misaligned with the more realistic scenarios like social media where opinions are rich in digital contexts. By design, surveys strip away the social and cultural context that shapes public opinion, and LLM benchmarks built on this paradigm inherit these critical limitations. To bridge this gap, we introduce MindVote, the first benchmark for public opinion prediction grounded in authentic social media discourse. MindVote is constructed from 3,918 naturalistic polls sourced from Reddit and Weibo, spanning 23 topics and enriched with detailed annotations for platform and topical context. Using this benchmark, we conduct a comprehensive evaluation of 15 LLMs, revealing a critical “survey-based specialization pitfall” where models fine-tuned on traditional surveys underperform their general-purpose counterparts and demonstrating the necessity of context in social media. MindVote provides a robust, ecologically valid framework to move beyond survey-based evaluations and advance the development of more socially intelligent AI systems.
@@ 1 Introduction
A core application for Large Language Models (LLMs) in understanding public opinion is serving as a rapid, scalable tool for pilot testing—predicting how a population will respond to a question before deploying costly, large-scale surveys (Rothschild et al. 2024; Bisbee et al. 2024; Suh et al. 2025; Sinacola, Pachot, and Petit 2025; Cao et al. 2025a; Anthis et al. 2025; Binz et al. 2025). This capability is crucial across a wide spectrum of domains. For example, LLMs can be used to predict reactions to new entertainment content and technological innovations, understand career aspirations, navigate complex social issues, and track evolving lifestyle trends (Argyle et al. 2023; Qian et al. 2025; Chen, Wang, and Ma 2025). However, the central goal is not simply to identify the majority opinion, but to predict how opinions are distributed across all possible choices. Understanding the full distribution is important because it reveals whether society is unified, divided, or contains strong dissent—information that a single majority number cannot capture. (Zhou, Nie, and Bansal 2022; Moon et al. 2024; Meister, Guestrin, and Hashimoto 2025; Cao et al. 2025b).
@@
However, the prevailing paradigm for public opinion distribution prediction relies on traditional structured surveys (Krogstad, Passel, and Cohn 2016; Santurkar et al. 2023; Elkjær and Wlezien 2024; Suh et al. 2025). This approach does not align with the fast-paced, context-rich digital environments—such as social media platforms—where opinions are now actively formed and expressed (Boutyline and Soter 2021; Riva, Aureli, and Silvestrini 2022). While surveys include basic demographic questions—such as age, gender, or education level—these static variables are poor proxies for the dynamic, situated nature of social context (Zaller and Feldman 1992). Demographics may indicate who a person is in broad categories, but fail to reflect how opinions are specifically shaped by lived experiences, group affiliations, ongoing conversations, or culturally reasoning patterns (Turner, Brown, and Tajfel 1979; Kahan, Jenkins-Smith, and Braman 2011; Markus and Kitayama 2014; McIntosh and Wright 2019).
@@
Consequently, LLM benchmarks built on this flawed paradigm, such as OpinionQA and SubPop (Santurkar et al. 2023; Suh et al. 2025), inherit these critical limitations. This reveals a crucial gap in the responsible deployment of LLMs. As these models are increasingly used to act as a proxy for human populations, ensuring their accuracy becomes essential. A proper evaluation framework must identify when models fail at context-rich, real-world tasks. Without this, we risk promoting systems that misrepresent the groups they are meant to simulate. This misrepresentation has severe downstream consequences, potentially compromising strategic decisions across critical topics like public policy, marketing, technology development, and community management (Qi 2024; Radivojevic, Clark, and Brenner 2024; Qu and Wang 2024; Panickssery, Bowman, and Feng 2024; Alekseenko et al. 2025; Wang, Morgenstern, and Dickerson 2025). This reliance on traditional survey-based evaluation creates three critical gaps in our ability to assess an LLM’s true social understanding:
@@
The Gap about Topic Imbalance. Real-world online discourse is thematically diverse (Tran and Ostendorf 2016; Song et al. 2023); for instance, entertainment drives 70% of traffic on Weibo (China Marketing Corp 2024), while Reddit’s largest communities focus on gaming and technology (Proferes et al. 2021). Yet, current survey-based benchmarks are heavily skewed towards formal, institutional topics like politics, which represent only a fraction of naturalistic opinion expression (Santurkar et al. 2023; Zhao et al. 2024). Because of this mismatch, we cannot evaluate if a model can adjust its reasoning from formal political surveys to the everyday language used on social media (Karjus and Cuskley 2024; Reveilhac and Morselli 2024).
@@
The Gap about Cultural Homogeneity. Existing cross-cultural survey benchmarks often exhibit cultural homogeneity, creating test conditions on Western-centric questions even with other languages translation (Haerpfer et al. 2022; DURMUS et al. 2024; Zhao et al. 2024). A model might perform well on a translated question, but we are left unable to determine if this success stems from genuine cultural understanding or a superficial linguistic competence (Singh et al. 2025).
@@
The Gap about Context Poverty. Opinions on social media are not noise but are helpful predictive signals (López-Rabadán 2021). This is correlated with context priming, a phenomenon where exposure to a specific context influences how individuals perceive and respond (Doyle and Lee 2016). On social platforms, users are primed by factors such as platform norms (e.g., Reddit’s anonymity vs. Weibo’s public-facing profiles (Brown et al. 2018; Zhu 2024)), temporal events (Banducci and Stevens 2015), and community-specific discourse (Tan et al. 2016; Qiao 2023). In contrast, survey-based benchmarks systematically remove this contextual priming to achieve standardization (Tourangeau, Singer, and Presser 2003), which prevents a true evaluation of a model’s ability to leverage these critical predictive cues (Li et al. 2023).
@@
To overcome these limitations, we move beyond surveys and instead use naturalistic social media polls—a more direct source for analyzing public opinion formation (Scarano et al. 2024). These polls address three critical gaps and provide a robust foundation for evaluation. The comparison between existing survey-based benchmarks and our benchmark is in Figure 1.
@@
Our Contributions. We introduce MindVote, the first benchmark for public opinion distribution prediction grounded in realistic social media discourse. We advance the field with the following innovations:
  • We construct and release MindVote, a benchmark of 3,918 authentic polls from two platforms (Reddit and Weibo) in their native English and Chinese. It spans 5 major topics and 23 sub-topics (shown in Figure 2) and is enriched with detailed social context annotations.
  • We benchmark leading LLMs, identify top performers, and reveal that models fine-tuned on traditional surveys face a critical survey-based specialization pitfall.
  • We demonstrate that enhancing a model’s capacity with social-context reasoning is more effective than fine-tuning on context-stripped data, offering a new direction for developing more socially intelligent systems.
@@ 2 Related Work
Opinion Distribution Prediction Benchmarks. U.S.-focused benchmarks like OpinionQA (Santurkar et al. 2023) and SubPop (Suh et al. 2025) concentrate on political topics, while international efforts like GlobalOpinionQA (DURMUS et al. 2024) is Western-centric frameworks and WorldValuesBench (Zhao et al. 2024) is framed as multi-cultural value prediction. Cao et al. (2025a) introduces three datasets (two English, one Chinese) specifically for group-level distribution prediction. These approaches contain only demographic-value pairings, structured survey format, or Western-centric lens that abstract away the naturalistic cultural contexts that are essential for authentic opinion distribution prediction.
@@ 3.1 Dataset Construction
Multi-Platform Data Collection. We collected 7,648 polls across two platforms spanning 2019-2025. Weibo dataset contributed 3,757 polls (2,026 from existing datasets (Lu et al. 2021) from 2019 to 2021, 1,734 newly crawled) and are anonymized for users’ personal identifiable information, capturing Chinese social media dynamics across pandemic and post-pandemic periods. Reddit provided 3,891 polls from diverse subreddits including r/poll during 2021-2025.
@@
    Content Quality Filtering: This filtering begins with duplicate content removal using a MinHash algorithm to eliminate items with greater than 95% overlap (Shrivastava and Li 2014), effectively targeting redundant trending topics, reposted polls, and cross-platform duplicates. Subsequently, automated toxicity screening was conducted using the Google Perspective API, with polls retaining a toxicity score below 0.4 being retained to filter out hate speech and overly controversial subjects (Lees et al. 2022).
@@
    Social Relevance Validation: Human verification of voting patterns (removing polls $\leq$ 100 votes for social relevance) and verification of poll and vote existence in all two platforms to ensure authentic social engagement and meaningful community participation.
@@
We manually annotate each poll includes rich metadata for social context enrichment: general platform context (e.g., user statistics and user behaviors) and topical context (topic-specific discourse patterns).
Cross-Cultural Augmentation. We created a parallel bilingual corpus for all two platforms through machine translation with rigorous quality control to demonstrate both linguistic and cultural effects: 10% back-translation validation (BLEU >> 35 (Papineni et al. 2002)), 5% native speaker rewriting, and expert review (Fleiss’ $\kappa$ = 0.51). The total number of polls becomes 3,918 after augmentation.
Final Dataset Composition. The final MindVote dataset is composed of 3,918 polls, with 2,158 sourced from Reddit and 1,760 from Weibo. Each poll is enriched with a comprehensive set of metadata, including its creation time, total vote count, and layers of social context: general platform context and topical context within that platform. To ensure broad accessibility and ease of use, the entire dataset is provided in both CSV and JSON formats. Table 1 shows an example with its simplified metadata keywords.
@@ 3.2 Experiment Design
Models. We evaluate 15 leading LLMs across closed-source, open-source and specialization categories: Claude-3.7-Sonnet-02-24 (Anthropic 2025), GPT-4o-2024-11-20 (OpenAI 2025b), GPT-4.1-04-14 (OpenAI 2025a), Gemini-2.5-Pro-05-06 (Google Cloud 2025), o3-medium-04-16 (OpenAI 2025c), Deepseek-R1-05-28 (DeepSeek-AI et al. 2025), Qwen2.5-32B-Instruct (Qwen Team 2024), Gemma-2-9B-it (Team, Gemma et al. 2024), Mistral-7B-Instruct-v0.1 (Chaplot 2023), Llama-2-13B-Chat (Touvron et al. 2023), Llama-3-70B-Instruct (Grattafiori et al. 2024), and Llama-4-Maverick-17B-128E (Meta AI 2025). We also evaluate three opinion distribution prediction specialization models from Suh et al. (2025). These models are fine-tuned using LoRA (Hu et al. 2022) on base models: Llama-3-70B-Base (Grattafiori et al. 2024), Llama-2-13B-Base (Touvron et al. 2023), and Mistral-7B-v0.1 (Chaplot 2023) which we named them respectively as: SubPop-Llama-3-70B, SubPop-Llama-2-13B, SubPop-Mistral-7B.
@@
Pipeline. All primary evaluations use a greedy decoding strategy (temperature=0) with default hyperparameter settings under zero-shot with context annotation (Kojima et al. 2022; Han et al. 2025). For the specialization fine-tuned models, we adapt those models into our pipeline by first loading the respective pretrained base model and then applying the publicly available LoRA weight checkpoints provided by Suh et al. (2025).
@@
Prompt. To ensure consistent and machine-readable outputs, we employ a structured prompting strategy where the model is given a JSON object containing the poll and its context. The template instructs the model to assume the role of a “opinion distribution prediction expert analyzing voting patterns and social dynamics.” The prompt includes the poll question and is enriched with social context metadata, with instruction for step-by-step reasoning. The model’s task is to return a JSON object with a schema identical to the input, replacing placeholder fields with its numeric predictions for the voting distribution.
@@
Evaluation Metrics. We adopt four distinct metrics to provide a comprehensive evaluation. Our primary metric is 1 - Wasserstein Distance (1-Wass.) (Santurkar et al. 2023; Meister, Guestrin, and Hashimoto 2025; Suh et al. 2025). The Wasserstein Distance measures the minimum cost for transforming one distribution into another, crucially accounting for semantic similarity between answer choices by treating them as points in a metric space. To complement this, we also report Spearman’s Rank Correlation Coefficient ($\rho$) (Zhou, Nie, and Bansal 2022; Pavlopoulos and Likas 2023), a non-parametric measure of how well the predicted ranking of options matches the true ranking of vote shares; 1 - KL Divergence (Meister, Guestrin, and Hashimoto 2025; Nguyen et al. 2025), which quantifies the information loss when using the model’s predicted distribution to approximate the ground truth; and One-hot Accuracy (Zhou, Nie, and Bansal 2022; Santurkar et al. 2023; Suh et al. 2025), which provides a strict measure of whether the single most likely predicted answer is correct.
@@
Evaluation Boundary. We include upper bounds and lower bounds for comparisons following (Suh et al. 2025). The upper bound is established by sampling subsets of the original results, calculating the four metrics between subsampled and original distributions, and performing bootstrapping to obtain a robust estimate that captures the intrinsic variance arising from the respondent sampling process in opinion. The uniform distribution lower bound establishes a performance floor equivalent to random chance.
4 Results
4.1 Performance of General Purpose Models
4.2 Survey-based Specialization Pitfalls
Our results in Table 2 show that fine-tuning on a structured survey dataset leads to a significant decrease in performance when predicting opinion distributions on social media. To demonstrate this, we compare our specialization models (base models that fine-tuned on traditional but context-stripped survey data) against their original base instruction-tuned models (the general-purpose models).
As shown in Table 2, specialization models consistently underperform their base counterparts, even when evaluated with full context annotations. For example, the SubPop-Llama-2-13B specialization model experiences a 3.3 percentage point decrease in its 1-Wass. score relative to the base Llama-2-13B. These results indicate that fine-tuning on sanitized data does not equip models to handle the complexities of real-world contexts. Rather than enhancing performance, such narrow specialization actually impairs the models’ ability to generalize to authentic social discourse.